# On complexity of alternating link equivalence

### v1

@@ Abstract.
Link equivalence up to isotopy in a 3-space is the problem that lies at the root of knot theory, and is important in 3-dimensional topology and geometry. We consider its restriction to alternating links, given by two alternating diagrams with $n_{1}$ and $n_{2}$ crossings, and show that this problem has polynomial algorithm in terms of $max\{n_{1},n_{2}\}$. For the proof, we use Tait flyping conjectures, observations from the work of Lackenby, Menasco, Sundberg and Thistlethwaite on alternating links, and algorithmic complexity of some problems from graph theory and topological graph theory.
@@ 1. Introduction
Links are important in the study of 3-manifolds, which is reflected, for example, in two classical theorems: Gordon-Luecke [13] and Lickorish-Wallace [26, 27, 43]. In recent years, interest in the algorithmic complexity of problems from 3-dimensional topology and geometry has been growing. Some problems are now known to be decidable or in NP (e.g. [25, 36, 21]), while some problems were proved to be NP-hard ([1, 20, 19, 10, 11, 12, 24]).
In 1960’s and 1970’s, link equivalence was proven to be decidable by Haken and Hemion [16, 15] using theory of normal surfaces. More recent proof, from 2011, is by Coward and Lackenby [8], and uses Reidemeister moves. Such diagrammatic moves were proved to connect any two diagrams of the same link in 1920’s by Alexander, Briggs and Reidemeister [2, 34]. Another proof that link equivalence is decidable can be obtained from the problem of 3-manifold homeomorphism, by passing from a link diagram to the triangulation of the link complement in 3-sphere. As detailed in the work of Kuperberg [21], the 3-manifold homeomorphism problem is known to be decidable due to the proof of Geometrization (conjectured and partly proven by Thurston [41], and completed by Perelman based on the program of Hamilton [32]). While links are not uniquely determined by their complements, one can check whether there is a homeomorphism between the link exteriors taking meridians to meridians (see Corollary 6.1.4 in Matveev’s book [29]), and this is enough for link isotopy. There is also an alternative solution to the homeomorphism problem for hyperbolic link complements by Dahmani and Groves [9], based on work of Sela [37] on isomorphism problem for hyperbolic groups.
Despite such a long and prolific history of study, no lower bounds on complexity of link equivalence are known. Is it NP-hard? Is there a polynomial algorithm? Here, we address its restriction to one infinite family, alternating links, and prove that this restriction is in class P (also called PTIME). Here is the restricted decision problem:
@@
Alternating links have been an important class of links: for example, they can be characterized purely topologically [17, 14], and their topological and geometric properties can often be seen or computed directly from a minimal crossing link diagram [30, 22, 39].
Over a century ago, Tait proposed the following conjecture: given prime alternating reduced diagrams $D_{1}$, $D_{2}$ of a link, one can transform $D_{1}$ into $D_{2}$ (up to planar isotopy) by a sequence of diagrammatic moves called flypes. A flype is depicted in Figure 1. The shaded disc labelled $T$ represents a 2-tangle in the terminology of Conway (i.e. a tangle with four free ends). In performing a flype, $T$ is turned by 180 degrees so that the crossing to its left is removed by twisting this tangle, while a new crossing is created to its right. The conjecture was proved in full generality by Menasco and Thistlethwaite as Theorem 1 in [31].
@@
We start our discussion of complexity of alternating link equivalence by showing in Section 3 that there are at most $2n^{2}$ flypes between two prime, alternating, reduced and connected diagrams of a link $L$ with $n$ crossings. We then proceed to give an algorithm in Section 4 that compares two such diagrams to determine if they belong to the same link. In Section 5, we check that the number of steps is polynomial in $n$ by reducing certain steps to problems on graphs. This implies that the complexity is polynomial for any two alternating diagrams (without connected, prime, reduced assumptions) due to topological results by Menasco [30].
The fact that alternating link equivalence is polynomial was conjectured, for example, in a recent paper by Samperton [35], as well as in some earlier talks and discussions on knots. The proof is based on known algorithmic complexity of problems from graph theory and topological graphs theory. Other facts that we use come from Tait flyping theorem and Lackenby’s work on volume of alternating links [22] from 2004, though some of these observations were likely known earlier: for example, some appear in the work of Sundberg and Thistlethwaite from 1998 [38].
Note that our algorithm does not immediately apply to alternating links that are not given by alternating diagrams. If a non-alternating diagram is given, there is an algorithm that determines whether the link is alternating given by Howie [17]. If it is, the algorithm produces its alternating diagram up to chirality. The complexity of this algorithm is exponential, since it uses normal surface theory.
@@ 2. Preliminaries: alternating links, their diagrams, and graphs
Let $L$ be an alternating link, i.e. such that its minimal crossing diagram is alternating. In an alternating diagram, overpasses and underpasses alternate as one travels along the link. Recall that a link diagram is prime if any simple closed curve that intersects the diagram in exactly two points has only an unknotted arc on one of its sides. A diagram is reduced if it does not contain nugatory (i.e. reducible) crossings. Consider a connected prime alternating reduced diagram $D$ of $L$ with $n$ crossings. Due to Menasco’s results [30], the link $L$ must be topologically prime and non-split. Reduced alternating diagram must also have the minimal crossing number, say $n$, due to another Tait conjecture, proved by Kauffman, Murasugi, and Thistlethwaite [18, 33, 40].
@@
Let us recall the terminology from Section 4 of Lackenby’s paper [22]. A square is a simple closed curve that intersects a link diagram four times. If the intersections lie on four different edges of the link diagram, the square is essential. A square is characteristic if it is essential, does not separate off a single crossing on either side, and any other square can be isotoped off it, where the isotopy is planar (and hence one cannot isotop through the crossings of $D$). A characteristic collection of squares is defined as follows: take one isotopy class of each characteristic square, and isotop them so that they are all disjoint.
@@
For a collection $\mathcal{C}$ of disjoint essential squares, a product region is defined as follows. Color the essential squares and their interior in blue. Consider $D$ without the blue interior. The regions bounded by arcs of $D$ (edges) and blue arcs on the boundary of squares can be colored in black and white in checkerboard fashion, since diagram regions of any reduced alternating diagram can be colored this way. If for color X (X is black or white), all regions of color X intersect crossings of $D$ and blue squares in only two connected components (e.g. in two crossings, or in a crossing and a blue arc), call the union of all regions of color X a complementary product region for $\mathcal{C}$. Lackenby demonstrated in Lemma 8 of [22] that a product region consists of twists and tangles beaded on two strands of a link as in Figure 2. This construction was also considered earlier by Sundberg and Thistlethwaite [38], and called a flyping circuit there.
Figure 2. Essential squares (in blue) and their complementary regions. In this figure, one can see a product region.
A flype on any essential square of $\mathcal{C}$ will both flip the square and move it one crossing to the left or right along the product region. A sequence of flypes can exchange such a square with a twist, or a subtwist in the same product region, but cannot exchange the squares between themselves.
@@
[22, Lemma 9 and Lemma 10] Let $\mathcal{C}$ be a collection of disjoint non-parallel essential squares. It is a characteristic collection if and only if all of the following hold:
@@
The condition (2) in the above lemma means the following situation does not occur: a square of a characteristic collection has two complementary regions on different sides of it (inside and outside), both of which are product regions, and their union is again a product region.
@@
The isomorphism problem for colored grahs with bounded color valence is known to be solvable in polynomial time. For example, Prop. 4.5 of [3] assures that canonical form of a vertex colored graph with color valence at most $d$ can be computed in $|V|^{P(d)}$ steps, where $P(d)$ is a polynomial that depends only on $d$. The graph canonization problem is at least as computationally hard as the graph isomorphism problem, so vertex colored graph isomorphism with bounded color valence has computational complexity at most $|V|^{P(d)}$ as well. This can be compared with graph isomorphism, which is not known to lie in P (and thus the problem for colored graphs is potentially easier).
@@ 3. Bound on the number of flypes
Bounds on purely diagrammatic moves that take one link diagram to another one have been a rich topic of study. Previous work concerns mostly Reidemeister moves (e.g. [8, 23]), but it is perhaps useful to have a simple bound on the number of flypes as well. More importantly for us, the facts needed for the polynomial algorithm are established in the proof of the following proposition.
@@
If $k=1$ and there are no subtwists from $P$ in $T$, call a flype that involves the tangle $T$ simple. Otherwise call it a concatenation flype. A concatenation flype can be substituted by a sequence of simple flypes. Hence, in what follows, we may consider only simple flypes, and count them for an upper bound. Suppose we have a sequence of diagrams $D=D_{1},D_{2},D_{3},...,D_{m}=D^{\prime}$, where $D_{i+1}$ is obtained from $D_{i}$ by a simple flype. We need to bound $m$.
Once a flype is applied to a diagram, some squares of $\mathcal{C}$ might move with the flype along their product regions, but otherwise the squares do not change, and a characteristic collection is still a characteristic collection after a flype, in the resulting diagram. The proof of this is straightforward and can be found in the first paragraph of Lemma 4 in [22]. In particular, $|\mathcal{C}|=|\mathcal{C}^{\prime}|$.
Now let’s count possible flypes that take us from a diagram $D$ to a diagram $D^{\prime}$ of $L$. In the initial diagram $D$, we have up to $n$ squares of $\mathcal{C}$. Indeed, each square $c$ of $\mathcal{C}$ has more than one crossing, but $c$ might have some other squares of $\mathcal{C}$ inside it. So we assume that for a square $c$, there is at least one crossing $x$ such that $x$ is not contained in other squares of $\mathcal{C}$ that do not contain all of $c$. There are also up to $n$ product regions in $D$, since each such region should contain at least one essential square or crossing.
@@ 4. Algorithm for alternating link equivalence
In this section, we give a simple algorithm for alternating link equivalence. In this section, we assume that alternating link diagrams $D_{1},D_{2}$ are prime, reduced, and connected. Once we have an algorithm for such diagrams, the algorithm for any two alternating diagrams is discussed, for example, in [35], based on the work of Menasco [30].
@@ 4.1. Basis for the algorithm
As noted above, the crossing numbers of $D_{1}$ and $D_{2}$ must be equal if the links are isotopic.
The following facts follow from the definition of a flype or are explained in Proposition 2 proof.
@@
    Characteristic collection $\mathcal{C}_{i}$ of $D_{i}$ is invariant under flypes: its squares may move inside their product regions, but no squares are removed or added to $\mathcal{C}_{i}$ when a flype is performed.
@@
    The 2-tangles (i.e. their projections in a link diagram) inside the squares of $\mathcal{C}_{i}$ do not change beyond planar isotopy when flypes are performed.
@@
    Every non-trivial flype can be substituted by a finite sequence of simple flypes such that for each simple flype, there is a square of $\mathcal{C}_{i}$ encircling the tangle $T$ that will be flyped.
@@
    Flypes cannot exchange squares of $\mathcal{C}_{i}$ in a product region.
@@
    A product region $P$ is uniquely determined by tangles inside the squares of $\mathcal{C}_{i}$ that $P$ contains, the cyclic order and number of squares (this number is called weight in [38]), and by the number of crossings in $P$ outside of the squares of $\mathcal{C}_{i}$ (call them free crossings of $P$).
Note that two equivalent product regions might look different, since we do not pay attention to the number of crossings between each two tangles, and only compare the total weight of the regions. But this is easily adjusted by flypes.
Therefore, to compare $D_{1}$ with $D_{2}$, it is enough to compare the tangles or product regions encircled by the squares of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, and then to compare how these squares “fit together” to make a link diagram. They might “fit together” in either product region or some other structure. (While we will not be using this, such structures are considered in [38] under the name “templates”, with pictures and examples.)
Squares of characteristic collection $\mathcal{C}_{i}$ may encircle tangles that have other squares of $\mathcal{C}_{i}$ in them, non-parallel to the initial square. If a link diagram is on a plane, each square has two sides: inside and outside. We call a side of a square of $\mathcal{C}_{i}$ inner if there are no other squares of $\mathcal{C}_{i}$ in the tangle that it encircles on that side. If there is only one square in a characteristic collection, it has two inner sides. We start by comparing tangles on the inner sides of squares, coloring them, then comparing tangles and regions which are right outside of them and coloring them next, etc., until we finally obtain two colored pseudographs of bounded valence instead of the two link diagrams we started with.
4.2. Algorithm
@@
Step 1. If $n_{1}\neq n_{2}$, stop, the links are not isotopic. If $n_{1}=n_{2}$, go to step 2.
Step 2. Detect a characteristic collection of squares $\mathcal{C}_{i}$ in $D_{i},i=1,2$

### v2

# Polynomial algorithm for alternating link equivalence

@@ Abstract.
Link equivalence up to isotopy in a 3-space is the problem that lies at the root of knot theory, and is important in 3-dimensional topology and geometry. We consider its restriction to alternating links, given by two alternating diagrams with $n_{1}$ and $n_{2}$ crossings, and show that this problem has polynomial algorithm in terms of $max\{n_{1},n_{2}\}$. For the proof, we use Tait flyping conjectures, observations stemming from the work of Lackenby, Menasco, Sundberg and Thistlethwaite on alternating links, and algorithmic complexity of some problems from graph theory and topological graph theory.
@@ 1. Introduction
In recent years, interest in the algorithmic complexity of problems from 3-dimensional topology and knot theory has been growing. Some problems were proved to be computationally hard, i.e. NP-hard ([1, 22, 21, 11, 12, 13, 27]). At the same time, almost nothing is known about the existence of such computationally easy non-trivial problems, i.e. the existence of polynomial algorithms in knot theory and 3-manifold topology. Here, we address restriction of link equivalence to one infinite family, alternating links, and prove that this restriction is in class P (also called PTIME), when the input is two alternating link diagrams. Here is the restricted problem:
@@
In 1960’s and 1970’s, link equivalence was proven to be decidable by Haken and Hemion [16, 15] using theory of normal surfaces. More recent proof, from 2011, is by Coward and Lackenby [9], and uses Reidemeister moves. Such diagrammatic moves were proved to connect any two diagrams of the same link in 1920’s by Alexander, Briggs and Reidemeister [2, 34]. Another proof that link equivalence is decidable can be obtained from the problem of 3-manifold homeomorphism, by passing from a link diagram to the triangulation of the link complement in 3-sphere. As detailed in the work of Kuperberg [23], the 3-manifold homeomorphism problem is known to be decidable due to the proof of Geometrization (conjectured and partly proven by Thurston [41], and completed by Perelman based on the program of Hamilton [32]). While links are not uniquely determined by their complements, one can check whether there is a homeomorphism between the link exteriors taking meridians to meridians (see Corollary 6.1.4 in Matveev’s book [29]), and this is enough for link isotopy. There is also an alternative solution to the homeomorphism problem for hyperbolic link complements by Dahmani and Groves [10], based on work of Sela [36] on isomorphism problem for hyperbolic groups. Meanwhile, practical link tabulation, often using supercomputers, is underway [17, 4, 38].
Alternating links are an important class of links: for example, they can be characterized purely topologically [18, 14], and their topological and geometric properties can often be seen or computed directly from a minimal crossing link diagram [30, 24, 39]. Over a century ago, Tait proposed the following conjecture: given prime alternating reduced diagrams $D_{1}$, $D_{2}$ of a link, one can transform $D_{1}$ into $D_{2}$ (up to planar isotopy) by a sequence of diagrammatic moves called flypes. A flype is depicted in Figure 1. The shaded disc labelled $T$ represents a 2-tangle in the terminology of Conway (i.e. a tangle with four free ends). In performing a flype, $T$ is turned by 180 degrees so that the crossing to its left is removed by twisting this tangle, while a new crossing is created to its right. The conjecture was proved in full generality by Menasco and Thistlethwaite as Theorem 1 in [31]. We will further call it Tait flyping theorem.
@@
The conjecture that alternating link equivalence is polynomial has circulated for a while. It has also been known that an existence of a polynomial algorithm for connected prime alternating reduced diagrams implies the existence of such algorithm for all alternating diagrams as a corollary of 1981 Menasco’s PhD thesis and paper [30]. This was, for example, used for a rapid and simple algorithm that decides whether an alternating diagram represents the unknot in Section 7 of Lackenby’s 2017 lecture notes [26]. So the main open question was the complexity of link equivalence if links are given by connected prime alternating reduced diagrams. It was noticed independently by different authors that flypes enforce even more structure on such diagrams than the Tait flyping theorem states: this was captured in the proof of this theorem by Menasco and Thistlethwaite [31], but also in the observations about flype circuits in the 1998 work of Sundberg and Thistlethwaite on the rate of growth of alternating links [37], and about characteristic squares and product regions in 2004 Lackenby’s work on volume of alternating links [24]. It was suggested that this extra structure might be a key to the polynomial algorithm (for example, by Lackenby), and yet the algorithm was never written. The conjecture and some of the previosly known observations about non-prime, non-reduced diagrams were recorded in algorithmic context in 2023 paper by Samperton that concerned quantum invariants of knots [35].
In this paper, we give the main piece: the polynomial algorithm for links presented by connected prime alternating reduced diagrams. In addition to the extra structure imposed on a link diagram by flypes, we use known problems from graph theory and topological graph theory for steps of the algorithm. Note that link isotopy problem cannot be reduced to any variation of graph isomorphism problem: one link can have many diagrams that give rise to many graphs. Instead, we detect tangles that must be equivalent in certain sense due to the aforementioned extra structure, and use plane embeddings of graphs and ordered lists of colors to compare them inductively, in a certain hierarchy.
Note that our algorithm does not immediately apply to alternating links that are not given by alternating diagrams. If a non-alternating diagram is given, there is an algorithm that determines whether the link is alternating given by Howie [18]. If it is, the algorithm produces its alternating diagram up to chirality. The complexity of Howie’s algorithm is exponential, since it uses normal surface theory.
A restricted problem, equivalence for knots (i.e. links with just one connected component) with alternating diagrams that satisfy some additional restrictions, was considered in [20], and was shown to have low time and space complexity using different methods.
@@
In what follows, we start by recalling some basic facts from knot theory, graph theory and complexity theory in Section 3. We then show in Section 4 that there are at most $2n^{2}$ flypes between two prime, alternating, reduced and connected diagrams of a link $L$ with $n$ crossings (Proposition 2). In Section 5, we outline the topological facts necessary for the algorithm justification, and give the algorithm that compares two such diagrams to determine if they belong to the same link. We prove that the algorithm works exactly as it should in Theorem 7. In Section 6, we check that the number of steps in the algorithm is polynomial in $n$.
2. Acknowledgements
We thank Marc Lackenby and Arnaud De Mesmay for helpful discussions and correspondence. Among other things, Marc Lackenby pointed us to some of the related statements in his work [24] back in 2010’s, and Arnaud De Mesmay gave us some helpful background concerning algorithms for plane graphs. We also thank Sofia Dinets for the help with some figures. Tsvietkova was partially supported by NSF DMS-2005496 and NSF CAREER DMS-2142487.
@@ 3. Preliminaries: alternating links, their diagrams, and graphs
We think of link diagrams in $\mathbb{R}^{2}$ or $\mathbb{S}^{2}$ interchangeably, only with some language differences: e.g. one speaks of a side of a closed curve in $\mathbb{S}^{2}$ versus being inside or outside of it in $\mathbb{R}^{2}$.
Let $L$ be an alternating link. Consider a connected prime alternating reduced diagram $D$ of $L$ with $n$ crossings. In Sections 3-5, we will only work with such diagrams. Due to Menasco’s results [30], the link $L$ must be topologically prime and non-split. Reduced alternating diagram must also have the minimal crossing number, say $n$, due to another Tait conjecture, proved by Kauffman, Murasugi, and Thistlethwaite [19, 33, 40].
@@
A 2-tangle in the terminology of Conway is a tangle with four free ends. 2-tangles in 3-space project to 2-tangles in $\mathbb{R}^{2}$ or $\mathbb{S}^{2}$. For 2-tangles in 3-space, we use ambient isotopy rel boundary, not allowing tangle ends to move. At the same time, there is usually no distinction between the four tangle ends unless we label the ends (we often will). Planar isotopy of 2-tangles implies their ambient isotopy rel boundary.
@@
Let us recall the terminology from Section 4 of Lackenby’s paper [24].
A square is a simple closed curve that intersects a link diagram four times. If the intersections lie on four different edges of the link diagram, the square is essential. A square is characteristic if it is essential, does not separate off a single crossing on either side, and any other square can be isotoped of it, where the isotopy is planar (this means pull one square continuously of another, eliminating all their intersections; in the process any intersection of each square with $D$ must stay in the same diagram edge where it was, and no new intersections with $D$ can be introduced). Figure 2(a) gives an example of two squares that are not characteristic. A characteristic collection of squares is defined as follows: take one isotopy class of each characteristic square, and isotope them so that they are all disjoint.
Figure 2. (A): a twist from a link diagram (in black) and two squares that are not characteristic, since they cannot be isotoped of each other. (B): essential squares (in blue) and their complementary regions. The product region is colored in grey.
@@
The following definition is crucial for the algorithm. A version of it can be found in [24]; we rephrase it slightly.
Color a collection $\mathcal{C^{\prime}}$ of essential squares in blue. Color the tangle each square in $\mathcal{C^{\prime}}$ encircles on one of its sides in blue as well. Consider $D$ without blue tangles, and denote it by $F$. The regions bounded by arcs of $F$ (edges) and blue arcs on the boundary of squares can be colored in black and white in checkerboard fashion, since diagram regions of any reduced alternating diagram can be colored this way. If for color X (X is black or white), all regions of color X intersect crossings of $F$ and blue squares in only two connected components (e.g. in two crossings, or in a crossing and a blue arc), and if all regions of color X are pairwise connected through crossings or pairwise adjacent to the same blue square, call the union of all regions of color X a complementary product region for $\mathcal{C^{\prime}}$. Lackenby’s observation in Lemma 8 of [24] implies that a product region consists of twists and tangles beaded on two strands of a link as in Figure 2(b). This construction was also considered earlier by Sundberg and Thistlethwaite [37], and called a flyping circuit there.
The above definition is for diagrams in $\mathbb{S}^{2}$: in particular, a product region can lie on any side of a square. Figure 2(b) shows one black product region for the blue squares.
Two distinct product regions can be adjacent to each other as follows: a square of $\mathcal{C}$ might have two product regions on different sides of it (inside and outside, if $D$ is in $\mathbb{R}^{2}$), with the two regions being connected through some edges connected in $D$. For example, the link diagram in Figure 4 (A) has adjacent product regions.
@@
[24, Lemma 9 and Lemma 10] Let $\mathcal{C}$ be a collection of disjoint non-parallel essential squares. It is a characteristic collection if and only if all of the following hold:
@@
The isomorphism problem for colored graphs with bounded color valence is known to be solvable in polynomial time. For example, Prop. 4.5 of [3] assures that canonical form of a vertex colored graph with color valence at most $d$ can be computed in $|V|^{P(d)}$ steps, where $P(d)$ is a polynomial that depends only on $d$. The graph canonization problem is at least as computationally hard as the graph isomorphism problem, so vertex colored graph isomorphism with bounded color valence has computational complexity at most $|V|^{P(d)}$ as well. This can be compared with graph isomorphism, which is not known to lie in P (and thus the problem for colored graphs is potentially easier).
@@ 4. Bound on the number of flypes
Bounds on purely diagrammatic moves that take one link diagram to another one have been a rich topic of study. Previous work concerns mostly Reidemeister moves (e.g. [9, 25]), but it is perhaps useful to have a simple bound on the number of flypes as well. More importantly for us, we establish a relation between potential flypes and squares of characteristic collection in the proof of the following proposition. We will use this relation later for the polynomial algorithm.
@@
If $k=1$ and there are no subtwists from $P$ in $T$, call a flype that involves the tangle $T$ simple. Otherwise call it a concatenation flype. A concatenation flype can be substituted by a sequence of simple flypes. Hence, in what follows, we may consider only simple flypes, and count them for an upper bound. Suppose we have a sequence of diagrams $D=D_{1},D_{2},D_{3},...,D_{m}=D^{\prime}$, where $D_{i+1}$ is obtained from $D_{i}$ by a simple flype. We need to bound $m$. We already established above that the tangles in $D_{1}$ to which we can apply simple non-trivial flypes are in bijective correspondence with the squares of $\mathcal{C}$.
Once a flype is applied to a diagram, some squares of $\mathcal{C}$ might move with the flype along their product regions, but otherwise the squares do not change, and a characteristic collection is still a characteristic collection after a flype, in the resulting diagram. The proof of this is straightforward and can be found in the first paragraph of Lemma 4 in [24]. In particular, $|\mathcal{C}|=|\mathcal{C}^{\prime}|$. Therefore, the tangles to which we can apply simple non-trivial flypes are in bijective correspondence with the squares of $\mathcal{C}$ not just in $D_{1}$, but also in any $D_{i},i=1,...,m$.
Now let’s count possible flypes that take us from a diagram $D$ to a diagram $D^{\prime}$ of $L$. Place $D$ in $\mathbb{R}^{2}$. Each square $c$ of $\mathcal{C}$ has more than one crossing, but $c$ might have some other squares of $\mathcal{C}$ inside it. If $c^{\prime}$ is inside $c$, call $c^{\prime}$ a subsquare of $c$, and is these squares are not parallel, call $c^{\prime}$ a proper subsquare of $c$. We may assume that for the square $c$, there is at least one crossing $x$ such that $x$ is not contained in any of its proper subsquares that are in $\mathcal{C}$. Hence the initial diagram $D$, we have up to $n$ squares of $\mathcal{C}$. There are also up to $n$ product regions in $D$, since each such region should contain at least one essential square or crossing.
@@ 5.1. Basis for the algorithm
The purpose of this subsection is to gather observations about alternating diagrams that yield a proof that our algorithm works. Some of these observations might be not new: implicitly, some of them seem to appear in [31, 37, 24]. However these previous papers address questions different from ours, and the exact statement of Lemma 4 is not there.
The following facts either follow from the definition of a flype or are explained in Proposition 2 proof.
@@
    Characteristic collection $\mathcal{C}$ of $D$ is invariant under flypes: its squares may move, but no squares are removed or added to $\mathcal{C}$ when a flype is performed.
@@
    A 2-tangle $T$ that lies inside a square $c$ of $\mathcal{C}$ does not change beyond planar isotopy of the tangle when a flype moving $c$ in $D$ is performed. (Note that $T$ is rotated with respect to $D$ after the flype.)
@@
    Every non-trivial flype can be substituted by a finite sequence of non-trivial simple flypes. The tangles to which simple flypes can be applied are in bijective correspondence with the squares of $\mathcal{C}$ (the squares encircle the respective tangles). A simple flype moves a tangle and respective square one crossing forward or backward in its product region $P$.
@@
    In particular, flypes cannot exchange squares of $\mathcal{C}$ in a product region.
We call a side of a square of $\mathcal{C}$ inner if it has no proper subsquares from $\mathcal{C}$ on that side. If there is only one square in a characteristic collection, it has two inner sides.
For a product region $P$, call crossings it contains that do not lie in squares of $\mathcal{C}$, free crossings of $P$.
From now on, we will place link diagrams $D,D^{\prime}$ in $\mathbb{R}^{2}$ for convenience.
Consider an extra curve that encircles all of $D$. Even though such a curve does not intersect $D$, we call it an exceptional square.
Remark 3.
Suppose we have two diagrams $D,D^{\prime}$, with one essential square each. Suppose two tangles inside these squares are planar isotopic, and two tangles outside the squares are planar isotopic as well. This does not mean that $D$ and $D^{\prime}$ are planar isotopic: for example, $D$ and $D^{\prime}$ and the links they represent can be related by a mutation, where one of the tangles in $D$ is rotated to obtain $D^{\prime}$. Therefore, it is important how we connect the tangles in $D$ and $D^{\prime}$. We will address this by labelling free ends of 2-tangles as follows: choose one free end if a 2-tangle and label it N, then proceed with labels W, S, E for the tangle ends counterclockwise. Assume the tangle is encircled by a square $c$: we label the ends twice, inside and right outside of the square. If such a tangle $T$ is isotopic to yet unlabelled tangle $T^{\prime}$, transfer the labels by the isotopy, and put them both inside and outside of the square encircling $T$. Now suppose $T_{1},T_{2}$ are 2-tangles lying in the same product region $P$ of $D$, with $k$ free crossings of $P$ between them, and the same is true for $T_{1}^{\prime},T_{2}^{\prime},P^{\prime},D^{\prime},k^{\prime}$. Suppose $T_{1}$ is planar isotopic to $T_{1}^{\prime}$, and $T_{2}$ to $T_{2}^{\prime}$. If $k<k^{\prime}$, we can perform $k^{\prime}-k$ flypes on $T_{2}^{\prime}$ in $P^{\prime}$ to make the number of free crossings between $T_{1}^{\prime}$ and $T_{2}^{\prime}$ equal to $k$. The flypes will rotate $T_{2}^{\prime}$ for $180(k^{\prime}-k)$ degrees. This might affect the labels at free ends of $T_{2}^{\prime}$, but it is simple to compare the labels of $T_{2},T_{2}^{\prime}$ taking this into account.
Lemma 4.
Suppose $C_{k+1}$ is either an exceptional square, or lies in a characteristic collection $\mathcal{C}$ for a diagram $D$. Consider inner squares $C_{1},...,C_{k}$ inside $C_{k+1}$. Take the fragment of $D$ that is all of $D$ inside $C_{k+1}$ and outside of $C_{1},...,C_{k}$. Denote this fragment by $F$. Then the following holds:
@@
    Either all crossings of $F$ belong to one product region $P$ that is a common product region for $C_{1},...,C_{k}$, or there is no product region in $F$.
  2. (2)
    Any simple non-trivial flype inside $C_{k+1}$ either moves one of $C_{i},i=1,...,k$, or happens fully within one of $C_{i}$.
  3. (3)
    Define $C_{1}^{\prime},...,C_{k+1}^{\prime},F^{\prime}$ for $D^{\prime}$ in the same way we defined $C_{1},...,C_{k+1},F$ for $D$. Suppose all crossings of $F,F^{\prime}$ belong to product regions $P,P^{\prime}$ respectively. Suppose further that the number of free crossings of $P,P^{\prime}$ is the same, the tangles in the sequences $C_{1},...,C_{k}$ and $C_{1}^{\prime},...,C_{k}^{\prime}$ are pairwise planar isotopic, the cyclic order of isotopic tangles coincides in the two sequences, and the positions of N, W, S, E labels in $F,F^{\prime}$ outside of the squares encircling $C_{1},...,C_{k}$ and $C_{1}^{\prime},...,C_{k}^{\prime}$ are the same after the respective rotation, as in Remark 3. Then the fragment of $D$ encircled by $C_{k+1}$ is ambient isotopic (rel boundary, if it is a tangle) to the fragment of $D^{\prime}$ encircled by $C_{k+1}^{\prime}$.
Note that number $k$ for $P$ is called weight in [37].
Proof.
Item (1) follows from the definition of product region, definition of $F$, and Lemma 1(1). In particular, any product region is a sequence of twists. If there is a product region $P$ inside $C_{k+1}$, and a crossing inside $C_{k+1}$ that does not belong to $P$, then some part of $P$ is encircled by an essential square $c$. By Lemma 1(1), $c$ is then parallel to the square of $\mathcal{C}$ that fully lies in $C_{k+1}$. Then $P$ is not in $F$ by the definition of $F$.
Item (2) follows from Lemma 1(2), item (3) in the discussion above the lemma, and (1) of this lemma. In particular, flypes only happen in product regions by item (3) in the discussion above the lemma. Two adjacent product regions are never compatible by Lemma 1(2). If there are two distinct product regions inside $C_{k+1}$, then one of them must lie inside $C_{i},i=1,...,k$ due to (1) of this lemma.
Item (3) follows from items (2)-(4) in the discussion above this lemma, the definition of the product region, and the Tait flyping theorem. Indeed, ambient isotopy for our diagrams is equivalent to planar isotopy together with flypes. The only way to change $P$ by flypes is to apply some combination of simple flypes. Every simple flype moves a tangle bounded by $C_{i}$ forward or backward for one crossing in $P$ by (3) in the discussion above this lemma. In the process, the number of free crossings of $P$ stays the same, as do the tangles bounded by $C_{1},...,C_{k}$, and their cyclic order due to items (2) and (4) in the discussion above. ∎
Remark 5.
For two link diagrams $D,D^{\prime}$, put a vertex at every crossing, and make diagram edges into edges between vertices. We obtain two planar pseudographs $\Gamma,\Gamma^{\prime}$, embedded in plane. Planar isotopy of $D,D^{\prime}$ is equivalent to plane isotopy of $\Gamma,\Gamma^{\prime}$ (such an isotopy is discussed in Subsection 3.3) together with the equivalence of crossing information for $D,D^{\prime}$. Note that for alternating link diagrams, it is enough to compare over/under information for just one crossing of $D$ with such information for the respective crossing of $D^{\prime}$: this determines similar information for the other crossings. If instead of all of $D,D^{\prime}$ we have two tangles $T,T^{\prime}$ encircled by squares $c,c^{\prime}$, we can form the graphs similarly, but need to put an extra vertex at every free end of each tangle. We put such vertices on the intersection of $T$ and $c$ for convenience. See Figure 3. Planar isotopy of $T,T^{\prime}$ is still equivalent to plane isotopy of the respective plane pseudographs together with the equivalence of crossing information for $T,T^{\prime}$.
Figure 3. On the left, a 2-tangle (in black) encircled by a square (in blue). On the right, the respective pseudograph (in black).
5.2. Algorithm and its justification.
In this subsection, we outline the algorithm, and prove that it works exactly as expected in Theorem 7.
Algorithm 6.
@@
Output: Yes or No, answering the question whether $D_{1},D_{2}$ belong to the same link.
Step 1. If $n_{1}\neq n_{2}$, stop with output No.
Step 2. Detect a characteristic collection of squares $\mathcal{C}_{i}$ in $D_{i},i=1,2$.